### v1

Abstract
Electronic health records (EHRs) enable powerful clinical prediction models but typically offer only coarse, post hoc explanations that are difficult to use in patient-level decision making. To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III. We map ICD-9 codes to PrimeKGchandak2023primekg, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes. We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models on the resulting supervision corpus. On ten PrimeKG-mapped diseases and small training cohorts (400 and 1,000 cases), our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47. Models transfer zero-shot to the CRADLE cohort, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
Introduction
Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLM can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.zhong2024; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions. However, applying CoT in medicine remains challenging. High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability. Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data. We propose a knowledge graph (KG)-guided chain-of-thought framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations. Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation. Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o. Guided by these paths, we prompt an LLM to generate temporally consistent, KG-anchored CoT explanations of whether a disease will appear at the next visit, and we retain only those reasoning traces whose conclusions match the ground-truth label. We then fine-tune lightweight open-weight models on this KG-anchored CoT supervision corpus so that, at inference time, they take ICD-9 features together with disease-specific KG evidence as input and produce both predictions and clinician-style explanations. This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
We evaluate this framework on ten PrimeKG-mapped diseases in MIMIC-III across small patient cohorts (400 and 1,000 patients), where our KG-anchored CoT supervision enables LLMs to outperform strong classical baselines such as XGBoost and Random Forest, achieving AUROC of approximately 0.67 and 0.70 and macro-AUPR of approximately 0.40 and 0.47, respectively. Clinician evaluations further show that KG-guided CoT yields substantially clearer, more coherent, and more clinically sound reasoning across three dimensions: clarity & coherence, coverage & relevance, and correctness & soundness. Finally, our CoT-tuned models transfer effectively to the CRADLE cohort for forecasting cardiovascular events within one year after type 2 diabetes, improving accuracy from roughly 0.40–0.51 (untuned LLMs) to 0.72–0.77 without any retraining, and providing more temporally consistent and clinically plausible explanations. Together, these results demonstrate that KG-anchored CoT supervision enables small LLMs to learn data-efficient, clinically grounded prognostic reasoning that generalizes across cohorts.
Method
Preliminaries.
A knowledge graph (KG) is a structured representation of entities and their relationships, typically modeled as a graph $G=(V,E)$ where $V$ denotes the set of entities (nodes) and $E$ the set of typed edges encoding relations among them. In the biomedical setting, KGs organize curated knowledge about diseases, phenotypes, drugs, genes, and related concepts, and provide an explicit relational backbone that can be used to guide or constrain model reasoning. Since biomedical relations in our external knowledge source are represented as direction- and type-specific (e.g., disease $\rightarrow$ gene and gene $\rightarrow$ disease may be encoded as distinct edges), we treat the underlying biomedical KG as a heterogeneous directed multigraph. Specifically, we adopt PrimeKGchandak2023primekg as our external knowledge source for grounding visit-level prognostic reasoning. PrimeKG is a large-scale biomedical KG that integrates curated relationships among diseases, phenotypes, drugs, genes, and other biomedical entities into a unified graph. We use it as the substrate for (1) aligning structured EHR concepts to KG nodes, (2) identifying disease-relevant neighborhood regions, and (3) extracting reasoning paths that connect clinical features to target diseases.
Task and Data.
We adopt the visit-level next-phenotype prediction task on MIMIC-III described in Section Results, following the 25 phenotyping labels defined by Harutyunyan et al.harutyunyan2019multitask Each index visit at time $t$ is encoded as a binary ICD-9 vector $x_{t}\in\{0,1\}^{7{,}423}$. Among the 25 candidate phenotypes, we focus on a subset $\mathcal{D}$ of 10 diseases that can be reliably mapped into PrimeKG; these 10 labels are the ones used for KG-guided CoT construction and downstream evaluation. For each disease $d\in\mathcal{D}$, we aim to predict whether $d$ will be present at the next visit t+1t{+}1. All models operate on the full ICD-9 feature space, and KG-based variants additionally consume disease-specific evidence derived from PrimeKG as described below.
Figure 1: General pipeline for KG-guided CoT data generation.
KG Entity Mapping.
To incorporate structured biomedical knowledge, we align ICD-9 concepts from MIMIC-III to PrimeKG nodes through a three-stage mapping procedure. Let $E_{\text{ICD}}$ denote the set of unique ICD-9 descriptions, and let $V_{G}$ be the node set of PrimeKG. For each entity $e_{i}\in E_{\text{ICD}}$, we retrieve a similarity-ranked candidate set $S_{i}=\{s_{1},\dots,s_{K}\}\subset V_{G}$ obtained by encoding $e$ and all KG node labels using a text embedding model and ranking candidates by cosine similarity. The final mapped entity $\hat{e}$ is selected in three stages:
Stage 1 (Exact match). If the ICD-9 text of $e_{i}$ exactly matches a node label $s_{k}$ in $S_{i}$, we record $\tilde{e}=s_{k}$.
Stage 2 (Similarity match). If an exact match is not found and the top similarity score exceeds a predefined threshold $\tau$ (set to 0.85), we select the most similar entity:
$\tilde{e_{i}}=\arg\max_{s_{k}\in S_{k}}\cos(e_{i},s_{k})\quad\text{if}\quad\max_{s_{k}\in S_{i}}\cos(e_{i},s_{k})>\tau.$
Stage 3 (LLM-based filtering). All provisional mappings $\{\tilde{e}_{i}\}_{i=1}^{N}$ from Stages 1 and 2 are then passed to GPT-4o, which validates or revises them to ensure clinical correctness, yielding the final mapped entities
$\hat{e}_{i}=\mathrm{LLM}(e_{i},\tilde{e}_{i}\mid I_{\text{select}}).$
where $I_{\text{select}}$ is an instruction prompt that asks the LLM to confirm, correct, or reject the candidate mapping based on clinical and biomedical plausibility. This three-stage procedure produces a final set of 1,513 mapped ICD-9 feature (phenotype) nodes and 10 mapped disease nodes, which serve as anchors for downstream relevance mining and reasoning path construction.
KG Relevance Node and Path Mining.
Let $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ denote the mapped ICD-9 feature nodes and $\{\hat{e}_{d}\}_{d\in\mathcal{D}}$ denote the mapped disease nodes in PrimeKG (here n=1,513n{=}1{,}513 and $|\mathcal{D}|{=}10$). Each $\hat{e}_{d}$ corresponds to one of the target diseases, while the $\hat{e}_{i}^{f}$ represent potential risk factors, comorbidities, or intermediate biomedical concepts. Our goal is to derive, for each disease $d\in\mathcal{D}$, a compact set of KG reasoning paths linking a small subset of feature nodes to that disease. For disease-centric node relevance, we present the disease node $\hat{e}_{d}$ and the set of all mapped feature nodes $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ to GPT-4o and identify the most relevant features for predicting the presence of $\hat{e}_{d}$ at t+1t{+}1. This yields a disease-level relevance set:
$\mathcal{R}_{d}=\{\hat{e}_{i,d}^{k}\}_{i=1}^{K}=\mathrm{LLM}(\{\hat{e}_{i}^{f}\}_{i=1}^{n},\hat{e}_{d}\mid I_{\mathrm{node\_select}}),\quad K\leq 8\;(\text{we set }K{=}8).$
Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$. Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
$\tilde{\mathcal{P}}_{i,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{i,d}^{k},\hat{e}_{d},G,L\big),$
$\{P_{i,d}^{k}\}_{k=1}^{K}=\mathrm{LLM}\!\big(\tilde{\mathcal{P}}_{i,d},\hat{e}_{d}\mid I_{\mathrm{prune}}\big),\quad K\leq 5\;\;(\text{we set }K{=}5).$
where $\mathrm{shortest\_path}(\cdot,\cdot;G,L)$ denotes the set of all minimum-length paths between two nodes in the biomedical knowledge graph $G$ (PrimeKG), subject to a maximum hop bound $L$ (we use L=5L{=}5 in our experiments) to avoid overly long, weakly supported chains. Since multiple equal-length paths may exist, many of which are clinically irrelevant, weakly supported, or redundant, we apply GPT-4o under a path-selection instruction $I_{\mathrm{path\_select}}$ to obtain a disease-level path set. Aggregating pruned paths across all $\hat{e}_{i}^{f}\in\mathcal{R}_{d}$ yields
$\mathcal{P}_{d}=\mathrm{LLM}\!\big(\{P_{i,d}^{k}\},\hat{e}_{d}\,\big|\,I_{\mathrm{path\_select}}\big),$
a compact subgraph capturing mechanistic or epidemiologic links from relevant clinical concepts to disease $d$. These subgraphs serve as the KG-grounded evidence used in our subsequent CoT generation.
CoT Generation with KG-Guided Reasoning.
Utilizing the disease-specific KG reasoning paths $\mathcal{P}_{d}$ as guidance, we distill reliable knowledge from the off-the-shelf KG into our CoT data. To achieve this, we prompt the LLM to consider these paths together with KG-related evidence and the observed feature nodes, and to elaborate them into medically grounded CoT explanations of the prediction, represented as
$C_{t,d}$ $=\mathrm{LLM}\big(d,\;x_{t}^{+},\;\mathcal{R}_{d}^{+}(t),\;\mathcal{R}_{d}^{-}(t),\;\mathcal{P}_{d},\;y_{t+1}^{(d)}\,\big|\,I_{\mathrm{gen}}\big),$
$\mathcal{R}_{d}^{+}(t)$ $=\big\{\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}\;\big|\;x_{t}(\hat{e}_{i,d}^{k})=1\big\},\quad\mathcal{R}_{d}^{-}(t)=\mathcal{R}_{d}\setminus\mathcal{R}_{d}^{+}(t).$
Here $I_{\mathrm{gen}}$ is the CoT instruction prompt, and $x_{t}\in\{0,1\}^{D}$ is the binary ICD–9 feature vector for the index visit at time $t$, with $x_{t}^{+}=\{c\mid x_{t}(c)=1\}$ denoting the set of ICD–9 codes present at that visit. The term $y_{t+1}^{(d)}\in\{0,1\}$ is the ground-truth label indicating whether disease $d$ is present at the next visit t+1t{+}1. The sets $\mathcal{R}_{d}^{+}(t)$ and $\mathcal{R}_{d}^{-}(t)$ indicate which disease-relevant KG nodes mapped from PrimeKG are expressed and not expressed at time $t$, respectively. The condition $x_{t}(\hat{e}_{i,d}^{k})=1$ means that the ICD–9 code corresponding to KG node $\hat{e}_{i,d}^{k}$ is present in the feature vector $x_{t}$.
Filtering.
Each generated chain-of-thought $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$. We use this conclusion to decide whether to keep the example: $C_{t,d}$ is retained if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$. The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs constitutes a compact but high-quality KG-grounded supervision corpus. We fine-tune lightweight open-weight LLMs (e.g., LLaMA, Gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
Figure 2: Schematic and example of KG-guided CoT generation and filtering.
Results
Dataset:
We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the phenotyping prediction setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits. For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1. This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{D}$ with D=7,423D=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit. The original task is formulated as a multi-label phenotyping problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 phenotyping labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1). For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1. Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict. We perform visit-level splitting over the 12,353 labeled index visits, randomly reserving 10% of cases as a held-out test set. From the remaining pool, we construct training sets of size 400 and 1,000 index visits to study data efficiency, using the rest for model selection.
In addition to MIMIC-III, we evaluate zero-shot transfer on Project CRADLE (Emory Clinical Research Analytics Data Lake Environment), a de-identified EHR repository from Emory Healthcare. Following prior work,xu2023hypehr we focus on patients with type 2 diabetes and formulate a binary prediction task: whether a patient will experience a
Table 1: The 10 MIMIC-III diseases mapped into PrimeKG and used as prediction targets.
Mapped Disease Label Category
Acute myocardial infarction Cardiovascular
Chronic kidney disease Renal
Chronic obstructive pulmonary disease Respiratory
Conduction disorders Cardiovascular
Coronary atherosclerosis Cardiovascular
Diabetes mellitus (no complication) Metabolic
Essential hypertension Cardiovascular
Gastrointestinal hemorrhage Gastrointestinal
Pneumonia Infectious
Shock Critical
cardiovascular disease (CVD) event within one year of the initial diabetes diagnosis. CVD endpoints include coronary heart disease, congestive heart failure, myocardial infarction, and stroke, identified via ICD-9/ICD-10 codes. After applying standard inclusion/exclusion criteria from the original CRADLE studies, the resulting cohort contains 36,611 patients with 12,724 binary features.
Evaluation Metrics: Since both the MIMIC-III and CRADLE cohorts exhibit substantial class imbalance, we therefore report accuracy, area under the receiver operating characteristic curve (AUROC), area under the precision–recall curve (AUPR), and macro-F1 as our primary evaluation metrics. For accuracy and F1, we convert predicted probabilities to binary decisions using a fixed threshold of 0.5. Table 1 lists the 10 diseases used as prediction targets.
Experiment:
We conduct experiments under two training regimes: one using 400 index-visit cases and another using 1,000 index-visit cases. Unless otherwise noted, all models are evaluated using a fixed decision threshold of 0.5 on the predicted probabilities.
Table 2: Data-efficiency comparison on MIMIC-III visit-level disease prediction with 400 and 1000 labeled cases (10 diseases). Best metric in bold, second-best underlined.
[…]
LLaMA3-8B (orig.) 0.4709 0.5604 0.2773 0.3327 0.4709 0.5604 0.2773 0.3327
[…]
Gemma-7B (orig.) 0.7475 0.5152 0.2528 0.0169 0.7475 0.5152 0.2528 0.0169
[…]
SGD 0.7277 $\pm$ 0.0251 0.6342 $\pm$ 0.0211 0.3595 $\pm$ 0.0235 0.1183 $\pm$ 0.0474 0.6886 $\pm$ 0.0243 0.6563 $\pm$ 0.0115 0.3840 $\pm$ 0.0114 0.1508 $\pm$ 0.0176
[…]
SVM 0.7385 $\pm$ 0.0000 0.6468 $\pm$ 0.0095 0.3824 $\pm$ 0.0032 0.1246 $\pm$ 0.0000 0.7391 $\pm$ 0.0011 0.6873 $\pm$ 0.0042 0.4256 $\pm$ 0.0054 0.1275 $\pm$ 0.0058
[…]
RF 0.7999 $\pm$ 0.0030 0.6155 $\pm$ 0.0084 0.3687 $\pm$ 0.0053 0.1970 $\pm$ 0.0043 0.8107 $\pm$ 0.0024 0.6574 $\pm$ 0.0076 0.4024 $\pm$ 0.0049 0.2191 $\pm$ 0.0042
[…]
XGBoost 0.8018 $\pm$ 0.0037 0.6476 $\pm$ 0.0053 0.3938 $\pm$ 0.0085 0.2090 $\pm$ 0.0058 0.8159 $\pm$ 0.0018 0.6799 $\pm$ 0.0095 0.4309 $\pm$ 0.0060 0.3065 $\pm$ 0.0058
As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting. With only 400 labeled cases, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$). Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant. When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming strong classical baselines such as XGBoost and Random Forest. These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
To evaluate cross-cohort robustness, we apply the MIMIC-trained models directly to the CRADLE dataset (36,611 patients, 12,724 binary features) without any re-training or calibration. The prediction task is cardiovascular disease onset within one year of type 2 diabetes diagnosis.
Table 3: Zero-shot transfer performance on the CRADLE cohort (cardiovascular disease within one year after T2D diagnosis). Best metric in bold, second-best underlined.
[…]
Llama3-8b-it Original 0.4026 0.5170 0.2241 0.3446
[…]
Gemma7b Original 0.5102 0.5052 0.2198 0.3064
[…]
In Table 3, MIMIC-trained models are effectively transferred to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes. For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while maintaining similar AUROC ($\approx 0.52$–0.53) and AUPR ($\approx 0.25$). Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to roughly 0.72–0.77 and macro-F1 from 0.31 to around 0.48, again with only modest changes in AUROC and AUPR. Results suggest that the decision boundaries learned from the KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
Human Evaluation:
To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline. For each of 100 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace—one from the KG-guided model and one from the baseline. Without knowing which model produced which trace, clinicians performed pairwise preferences, selecting the better explanation (A or B) along three dimensions:
  • Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions.
  • Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoid unnecessary detail.
  • Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better. The annotation was performed by two human experts from different medical schools: an MD professor and a PhD student in biomedical informatics. In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the PhD student reviewing 55 cases.
Figure 3: Clinician preferences for reasoning quality.
On the 115 annotated cases, both clinicians preferred the KG-guided model (Model A) over the untuned baseline (Model B), with overall 96.5% of cases for Clarity & Coherence, 94.8% of cases for Coverage & Relevance, and 98.3% of cases for Correctness & Soundness. Qualitatively, clinicians reported that KG-guided CoT traces more often (i) organized evidence in a temporally consistent manner, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details. By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES even when only weak signals were present). While both models are sensitive to poor or incomplete input structure, the untuned baseline was more prone to repetitive differentials and illogical reasoning, especially on shorter or noisier inputs.
Discussion
In this study, we introduced a KG-guided chain-of-thought framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data. By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we created a compact supervision corpus that supports data-efficient fine-tuning of lightweight LLMs.
Our results indicate that KG-guided CoT improves data-efficient prognostic prediction. On MIMIC-III visit-level disease prediction, KG-guided CoT fine-tuning substantially improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and brings them in line with strong classical baselines. With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$, matching or exceeding logistic regression and XGBoost while also producing explicit reasoning traces; with 1,000 visits, performance further improves to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$. These gains suggest that KG-anchored CoT supervision acts as an effective “reasoning prior,” helping small LLMs learn clinically meaningful decision boundaries from limited data. Classical models remain competitive on AUROC, but KG-guided LLMs tend to close or surpass this gap while improving macro-AUPR and F1, especially for minority conditions. Practically, this points toward hybrid designs where well-calibrated classical or deep risk models provide scores, and KG-guided LLMs serve as a clinician-facing layer that turns those scores into structured, patient-specific, KG-grounded explanations.
We also observe promising cross-cohort generalization and decision-boundary transfer. Our zero-shot transfer from MIMIC-III to CRADLE suggests that KG-guided CoT yields reasonably robust decision boundaries: without any retraining, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from about 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain relatively stable. This pattern indicates that overall discrimination is preserved, but the effective decision thresholds and class balance improve in the target cohort.
Human evaluation provides complementary evidence that KG guidance improves the clinical usability of explanations. In 115 blinded case comparisons between our KG-guided LLaMA model and an untuned LLM baseline, the KG-guided traces were preferred on clarity and coherence, coverage and relevance, and correctness and soundness in the large majority of cases. Qualitative feedback from the MD professor and PhD student indicated that KG-guided traces more often (i) organized evidence in a temporally consistent way, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details. By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES) and sometimes reached correct labels via illogical or poorly justified reasoning, especially on shorter or noisier inputs. These observations suggest that anchoring CoT to KG paths, combined with label-consistency filtering, shapes both the style and the content of reasoning toward disease-specific, clinically grounded narratives.
Several limitations remain. Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision. ICD–9-to-KG mapping and relevance/path selection rely on GPT-4o and heuristic thresholds, which may introduce subtle errors despite label-consistency filtering. Human annotators also highlighted opportunities to improve the input representation itself: our current summaries sometimes blur temporal structure, such as exposure versus outcome windows, and care setting (inpatient versus outpatient), which likely caps attainable performance and can lead even strong models to produce brittle or poorly justified reasoning. Finally, the current models require KG access at inference time for KG-conditioned variants.
Looking ahead, our results point to several concrete directions for future work. Applying the framework to richer EHR views, such as notes, medications, labs, and additional KGs could support more nuanced, multimodal reasoning, while replacing simple shortest-path guidance with richer graph reasoning may further improve explanation quality. Clinician feedback also points to a complementary direction: improving input structuring by making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and including basic demographics to reduce tendency for the model to hallucinate or over-interpret limited evidence. In parallel, future work may introduce temporal or guideline-based constraints to reduce illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
Conclusion
In this paper, we introduced a KG-guided chain-of-thought framework that couples biomedical structure from PrimeKG with LLM-based reasoning over structured EHR data. By using KG paths as scaffolds for CoT generation and label-consistent filtering, we constructed a compact supervision corpus that enables small open-weight LLMs to achieve competitive or superior visit-level disease prediction on MIMIC-III under strict data constraints, while also producing explicit, clinically grounded explanations. Zero-shot transfer to the CRADLE cohort suggests that the learned decision boundaries generalize across cohorts and feature spaces, and blinded clinician review indicates a consistent preference for KG-guided CoT over untuned LLM explanations. Taken together, these results highlight KG-anchored CoT supervision as a promising direction for building clinical decision-support systems in which LLMs not only predict risk but also articulate transparent, patient-level reasoning aligned with biomedical knowledge and clinician expectations.
Acknowledgements
This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University.
Appendix
Implementation Details
All fine-tuning experiments were conducted on a single NVIDIA H200 GPU on our institutional high-performance computing cluster, with 512 GB of host memory allocated via SLURM. We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT supervision corpus using full-parameter updates with AdamW (learning rate $1\times 10^{-5}$, cosine learning-rate schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096 tokens, and 10 training epochs in bf16 precision under DeepSpeed ZeRO-3. We monitored validation performance and selected the checkpoint with the best macro-AUPR. KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100.
Prompt used:
Figure 4: Prompt templates used in our KG-guided CoT pipeline: (top-left) disease-relevant node selection, (top-right) KG path pruning and selection, (bottom) CoT generation conditioned on KG evidence and visit features.

### v2

@@
-
+Affiliations withheld for double-blind review
 Abstract
@@
 To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III.
-We map ICD-9 codes to PrimeKGchandak2023primekg, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes.
-We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models on the resulting supervision corpus.
-On ten PrimeKG-mapped diseases and small training cohorts (400 and 1,000 cases), our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47.
-Models transfer zero-shot to the CRADLE cohort, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
+We map ICD-9 codes to PrimeKG, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes.
+We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models under two resulting training cohorts (400 and 1,000 index visits from MIMIC-III).
+On ten PrimeKG-mapped diseases and small training cohorts, our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47.
+Models transfer zero-shot to the CRADLE cohort without any additional fine-tuning, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
 Introduction
-Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
-Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLM can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.zhong2024; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions.
+Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation; zheng2014automatically As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.choi2016retain; ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
+Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.wei2022chain; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions.
 However, applying CoT in medicine remains challenging.
-High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability.
+High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.xu2024ram; liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability.
 Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
 In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data.
-We propose a knowledge graph (KG)-guided chain-of-thought framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations.
+We propose a KG-guided CoT framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations.
 Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation.
@@
 We use it as the substrate for (1) aligning structured EHR concepts to KG nodes, (2) identifying disease-relevant neighborhood regions, and (3) extracting reasoning paths that connect clinical features to target diseases.
-Task and Data.
-We adopt the visit-level next-phenotype prediction task on MIMIC-III described in Section Results, following the 25 phenotyping labels defined by Harutyunyan et al.harutyunyan2019multitask Each index visit at time $t$ is encoded as a binary ICD-9 vector $x_{t}\in\{0,1\}^{7{,}423}$.
-Among the 25 candidate phenotypes, we focus on a subset $\mathcal{D}$ of 10 diseases that can be reliably mapped into PrimeKG; these 10 labels are the ones used for KG-guided CoT construction and downstream evaluation.
-For each disease $d\in\mathcal{D}$, we aim to predict whether $d$ will be present at the next visit t+1t{+}1.
-All models operate on the full ICD-9 feature space, and KG-based variants additionally consume disease-specific evidence derived from PrimeKG as described below.
+Prediction Task.
+We consider a next visit disease prediction task over structured EHR data.
+For each patient, let $\{x_{t}\}_{t=1}^{T}$ denotes a sequence of visits ordered in time, where each visit is represented as a binary feature vector $x_{t}\in\{0,1\}^{N}$, with $N$ the dimensionality of the code space.
+Let $\mathcal{D}$ denotes a set of disease labels of interest, and for each $d\in\mathcal{D}$ and visit index $t$, define a binary outcome $y_{t+1}^{(d)}\in\{0,1\}$, indicating whether disease $d$ is present at the next visit t+1t{+}1.
+The model’s goal is, for each pair $(x_{t},d)$, to estimate a probability $\hat{p}^{(d)}_{t+1}=f(x_{t},d)$ and an associated binary decision $\hat{y}^{(d)}_{t+1}$ about whether $d$ will appear at visit t+1t{+}1.
+Our KG-guided models are additionally conditioned on disease-specific KG evidence (relevance sets and reasoning paths) associated with $d$, as described below.
+In our experiments, we instantiate this general setup using ICD–9 diagnosis codes as features, with N=7,423N=7{,}423 and $|\mathcal{D}|=10$ diseases on the MIMIC-III cohort (details in the Results section).
 Figure 1: General pipeline for KG-guided CoT data generation.
@@
 Let $E_{\text{ICD}}$ denote the set of unique ICD-9 descriptions, and let $V_{G}$ be the node set of PrimeKG.
-For each entity $e_{i}\in E_{\text{ICD}}$, we retrieve a similarity-ranked candidate set $S_{i}=\{s_{1},\dots,s_{K}\}\subset V_{G}$ obtained by encoding $e$ and all KG node labels using a text embedding model and ranking candidates by cosine similarity.
-The final mapped entity $\hat{e}$ is selected in three stages:
-Stage 1 (Exact match). If the ICD-9 text of $e_{i}$ exactly matches a node label $s_{k}$ in $S_{i}$, we record $\tilde{e}=s_{k}$.
+For each entity $e_{i}\in E_{\text{ICD}}$, we retrieve a similarity-ranked candidate set $S_{i}=\{s_{1},\dots,s_{C}\}\subset V_{G}$ obtained by encoding $e_{i}$ and all KG node labels using a text embedding model and ranking candidates by cosine similarity.
+The final mapped entity $\hat{e}_{i}$ is selected in three stages:
+Stage 1 (Exact match). If the ICD-9 text of $e_{i}$ exactly matches a node label $s_{c}$ in $S_{i}$, we record $\tilde{e}_{i}=s_{c}$.
 Stage 2 (Similarity match). If an exact match is not found and the top similarity score exceeds a predefined threshold $\tau$ (set to 0.85), we select the most similar entity:
-$\tilde{e_{i}}=\arg\max_{s_{k}\in S_{k}}\cos(e_{i},s_{k})\quad\text{if}\quad\max_{s_{k}\in S_{i}}\cos(e_{i},s_{k})>\tau.$
+$\tilde{e}_{i}=\arg\max_{s_{c}\in S_{i}}\cos(e_{i},s_{c})\quad\text{if}\quad\max_{s_{c}\in S_{i}}\cos(e_{i},s_{c})>\tau.$
 Stage 3 (LLM-based filtering).
 All provisional mappings $\{\tilde{e}_{i}\}_{i=1}^{n}$ from Stages 1 and 2 are then passed to GPT-4o, which validates or revises them to ensure clinical correctness, yielding the final mapped entities
-$\hat{e}_{i}=\mathrm{LLM}(e_{i},\tilde{e}_{i}\mid I_{\text{select}}).$
-where $I_{\text{select}}$ is an instruction prompt that asks the LLM to confirm, correct, or reject the candidate mapping based on clinical and biomedical plausibility.
-This three-stage procedure produces a final set of 1,513 mapped ICD-9 feature (phenotype) nodes and 10 mapped disease nodes, which serve as anchors for downstream relevance mining and reasoning path construction.
+$\hat{e}_{i}=\mathrm{LLM}(e_{i},\tilde{e}_{i}\mid I_{\text{select}}),$
+where $I_{\text{select}}$ is the prompt that asks the LLM to confirm, correct, or reject the candidate mapping based on clinical and biomedical plausibility.
+This three-stage procedure produces a final set of 1,513 mapped ICD-9 feature nodes and 10 mapped disease nodes, which serve as anchors for downstream relevance mining and reasoning path construction.
 KG Relevance Node and Path Mining.
@@
 For disease-centric node relevance, we present the disease node $\hat{e}_{d}$ and the set of all mapped feature nodes $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ to GPT-4o and identify the most relevant features for predicting the presence of $\hat{e}_{d}$ at t+1t{+}1.
-This yields a disease-level relevance set:
-$\mathcal{R}_{d}=\{\hat{e}_{i,d}^{k}\}_{i=1}^{K}=\mathrm{LLM}(\{\hat{e}_{i}^{f}\}_{i=1}^{n},\hat{e}_{d}\mid I_{\mathrm{node\_select}}),\quad K\leq 8\;(\text{we set }K{=}8).$
-Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$.
+This yields relevance set:
+$\mathcal{R}_{d}=\{\hat{e}_{k,d}\}_{k=1}^{K_{\text{node}}}=\mathrm{LLM}(\{\hat{e}_{i}^{f}\}_{i=1}^{n},\hat{e}_{d}\mid I_{\mathrm{node\_select}}),\quad K_{\text{node}}=8\;,$
+where each $\hat{e}_{k,d}$ is a mapped feature node selected from $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ for disease $d$.
+Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{k,d}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$.
 Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
-$\tilde{\mathcal{P}}_{i,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{i,d}^{k},\hat{e}_{d},G,L\big),$
-$\{P_{i,d}^{k}\}_{k=1}^{K}=\mathrm{LLM}\!\big(\tilde{\mathcal{P}}_{i,d},\hat{e}_{d}\mid I_{\mathrm{prune}}\big),\quad K\leq 5\;\;(\text{we set }K{=}5).$
+$\tilde{\mathcal{P}}_{k,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{k,d},\hat{e}_{d},G,L\big),$
+$\mathcal{P}_{d}=\{P_{k,d}\}_{k=1}^{K_{\text{path}}}=\mathrm{LLM}\!\left(\big\{\tilde{\mathcal{P}}_{k,d}:\hat{e}_{k,d}\in\mathcal{R}_{d}\big\},\hat{e}_{d}\,\middle|\,I_{\mathrm{path\_select}}\right),\quad K_{\text{path}}=5,$
 where $\mathrm{shortest\_path}(\cdot,\cdot;G,L)$ denotes the set of all minimum-length paths between two nodes in the biomedical knowledge graph $G$ (PrimeKG), subject to a maximum hop bound $L$ (we use L=5L{=}5 in our experiments) to avoid overly long, weakly supported chains.
-Since multiple equal-length paths may exist, many of which are clinically irrelevant, weakly supported, or redundant, we apply GPT-4o under a path-selection instruction $I_{\mathrm{path\_select}}$ to obtain a disease-level path set.
-Aggregating pruned paths across all $\hat{e}_{i}^{f}\in\mathcal{R}_{d}$ yields
-$\mathcal{P}_{d}=\mathrm{LLM}\!\big(\{P_{i,d}^{k}\},\hat{e}_{d}\,\big|\,I_{\mathrm{path\_select}}\big),$
-a compact subgraph capturing mechanistic or epidemiologic links from relevant clinical concepts to disease $d$. These subgraphs serve as the KG-grounded evidence used in our subsequent CoT generation.
+Since multiple paths may exist for a disease, many of which are clinically irrelevant, weakly supported, or redundant, we apply GPT-4o under a path-selection instruction $I_{\mathrm{path\_select}}$ to obtain a disease-level path set, yielding a compact subgraph capturing mechanistic or epidemiologic links from relevant clinical concepts to disease $d$.
+These subgraphs serve as the KG-grounded evidence used in our subsequent CoT generation.
 CoT Generation with KG-Guided Reasoning.
@@
 $C_{t,d}$ $=\mathrm{LLM}\big(d,\;x_{t}^{+},\;\mathcal{R}_{d}^{+}(t),\;\mathcal{R}_{d}^{-}(t),\;\mathcal{P}_{d},\;y_{t+1}^{(d)}\,\big|\,I_{\mathrm{gen}}\big),$
-$\mathcal{R}_{d}^{+}(t)$ $=\big\{\hat{e}_{i,d}^{k}\in\mathcal{R}_{d}\;\big|\;x_{t}(\hat{e}_{i,d}^{k})=1\big\},\quad\mathcal{R}_{d}^{-}(t)=\mathcal{R}_{d}\setminus\mathcal{R}_{d}^{+}(t).$
-Here $I_{\mathrm{gen}}$ is the CoT instruction prompt, and $x_{t}\in\{0,1\}^{D}$ is the binary ICD–9 feature vector for the index visit at time $t$, with $x_{t}^{+}=\{c\mid x_{t}(c)=1\}$ denoting the set of ICD–9 codes present at that visit.
+$\mathcal{R}_{d}^{+}(t)$ $=\big\{\hat{e}_{k,d}\in\mathcal{R}_{d}\;\big|\;x_{t}(\hat{e}_{k,d})=1\big\},\quad\mathcal{R}_{d}^{-}(t)=\mathcal{R}_{d}\setminus\mathcal{R}_{d}^{+}(t).$
+Here $I_{\mathrm{gen}}$ is the CoT instruction prompt, and $x_{t}\in\{0,1\}^{N}$ is the binary ICD–9 feature vector for the index visit at time $t$, with $x_{t}^{+}=\{c\mid x_{t}(c)=1\}$ denoting the set of ICD–9 codes present at that visit.
 The term $y_{t+1}^{(d)}\in\{0,1\}$ is the ground-truth label indicating whether disease $d$ is present at the next visit t+1t{+}1.
 The sets $\mathcal{R}_{d}^{+}(t)$ and $\mathcal{R}_{d}^{-}(t)$ indicate which disease-relevant KG nodes mapped from PrimeKG are expressed and not expressed at time $t$, respectively.
-The condition $x_{t}(\hat{e}_{i,d}^{k})=1$ means that the ICD–9 code corresponding to KG node $\hat{e}_{i,d}^{k}$ is present in the feature vector $x_{t}$.
+The condition $x_{t}(\hat{e}_{k,d})=1$ means that the ICD–9 code corresponding to KG node $\hat{e}_{k,d}$ is present in the feature vector $x_{t}$.
 Filtering.
-Each generated chain-of-thought $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$.
-We use this conclusion to decide whether to keep the example: $C_{t,d}$ is retained if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$.
-The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs constitutes a compact but high-quality KG-grounded supervision corpus.
-We fine-tune lightweight open-weight LLMs (e.g., LLaMA, Gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
+Each generated CoT $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$.
+We use this conclusion to decide whether to keep the example: $C_{t,d}$ is kept if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$.
+The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs forms a compact but high-quality KG-grounded supervision corpus.
+We fine-tune lightweight open-weight LLMs (e.g., LLaMA grattafiori2024llama, Gemma team2024gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
 Figure 2: Schematic and example of KG-guided CoT generation and filtering.
 Results
-Dataset:
-We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the phenotyping prediction setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits.
+Dataset.
+We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits.
 For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1.
 This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
-Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{D}$ with D=7,423D=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit.
-The original task is formulated as a multi-label phenotyping problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 phenotyping labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1).
+Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{N}$ with N=7,423N=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit.
+The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 disease labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1).
 For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1.
@@
 After applying standard inclusion/exclusion criteria from the original CRADLE studies, the resulting cohort contains 36,611 patients with 12,724 binary features.
-Evaluation Metrics: Since both the MIMIC-III and CRADLE cohorts exhibit substantial class imbalance, we therefore report accuracy, area under the receiver operating characteristic curve (AUROC), area under the precision–recall curve (AUPR), and macro-F1 as our primary evaluation metrics.
+Evaluation Metrics.
+Since both the MIMIC-III and CRADLE cohorts exhibit substantial class imbalance, we therefore report accuracy, area under the receiver operating characteristic curve (AUROC), area under the precision–recall curve (AUPR), and macro-F1 as our primary evaluation metrics.
 For accuracy and F1, we convert predicted probabilities to binary decisions using a fixed threshold of 0.5.
@@
 Unless otherwise noted, all models are evaluated using a fixed decision threshold of 0.5 on the predicted probabilities.
-Table 2: Data-efficiency comparison on MIMIC-III visit-level disease prediction with 400 and 1000 labeled cases (10 diseases). Best metric in bold, second-best underlined.
+Table 2: Data-efficiency comparison on MIMIC-III visit-level disease prediction with 400 and 1000 index visits (10 diseases). Best metric in bold, second-best underlined.
 […]
@@
 As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting.
-With only 400 labeled cases, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$).
+With only 400 labeled index visits data, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$).
 Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant.
@@
 […]
-Llama3-8b-it Original 0.4026 0.5170 0.2241 0.3446
+Llama3-8b-it (orig.) 0.4026 0.5170 0.2241 0.3446
 […]
 […]
-Gemma7b Original 0.5102 0.5052 0.2198 0.3064
+Gemma-7b (orig.) 0.5102 0.5052 0.2198 0.3064
 […]
 […]
-In Table 3, MIMIC-trained models are effectively transferred to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes.
-For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while maintaining similar AUROC ($\approx 0.52$–0.53) and AUPR ($\approx 0.25$).
-Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to roughly 0.72–0.77 and macro-F1 from 0.31 to around 0.48, again with only modest changes in AUROC and AUPR.
-Results suggest that the decision boundaries learned from the KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
-Human Evaluation:
+MIMIC-trained models transfer effectively to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes.
+For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while modestly raising AUROC (to $\approx 0.52$–0.53) and AUPR (to $\approx 0.25$).
+Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to 0.72–0.77 and macro-F1 from 0.31 to about 0.48, again with modest changes in AUROC and AUPR.
+These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
+Human Evaluation.
 To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline.
-For each of 100 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace—one from the KG-guided model and one from the baseline.
-Without knowing which model produced which trace, clinicians performed pairwise preferences, selecting the better explanation (A or B) along three dimensions:
+For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace, one from the KG-guided model and one from the baseline.
+Without knowing which model produced which trace, clinicians made pairwise preferences, selecting the better explanation (A or B) along three dimensions: 1.
 Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions.
-  • Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoid unnecessary detail.
+2.
+Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoids unnecessary detail.
+3.
 Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
 For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better.
-The annotation was performed by two human experts from different medical schools: an MD professor and a PhD student in biomedical informatics.
-In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the PhD student reviewing 55 cases.
+The annotation was performed by two human experts from different medical schools: an MD professor and a MD/PhD student in biomedical informatics.
+In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the MD/PhD student reviewing 50 cases.
 Figure 3: Clinician preferences for reasoning quality.
@@
 Discussion
-In this study, we introduced a KG-guided chain-of-thought framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data.
-By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we created a compact supervision corpus that supports data-efficient fine-tuning of lightweight LLMs.
-Our results indicate that KG-guided CoT improves data-efficient prognostic prediction.
-On MIMIC-III visit-level disease prediction, KG-guided CoT fine-tuning substantially improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and brings them in line with strong classical baselines.
-With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$, matching or exceeding logistic regression and XGBoost while also producing explicit reasoning traces; with 1,000 visits, performance further improves to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$.
-These gains suggest that KG-anchored CoT supervision acts as an effective “reasoning prior,” helping small LLMs learn clinically meaningful decision boundaries from limited data.
-Classical models remain competitive on AUROC, but KG-guided LLMs tend to close or surpass this gap while improving macro-AUPR and F1, especially for minority conditions.
-Practically, this points toward hybrid designs where well-calibrated classical or deep risk models provide scores, and KG-guided LLMs serve as a clinician-facing layer that turns those scores into structured, patient-specific, KG-grounded explanations.
-We also observe promising cross-cohort generalization and decision-boundary transfer.
-Our zero-shot transfer from MIMIC-III to CRADLE suggests that KG-guided CoT yields reasonably robust decision boundaries: without any retraining, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from about 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain relatively stable.
-This pattern indicates that overall discrimination is preserved, but the effective decision thresholds and class balance improve in the target cohort.
+In this study, we proposed a KG-guided chain-of-thought generation framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data.
+By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we construct a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs.
+On MIMIC-III visit-level disease prediction, KG-guided CoT improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and makes them competitive with strong classical baselines.
+With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$; with 1,000 visits, performance increases to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$.
+These results suggest that KG-anchored CoT supervision acts as an effective reasoning prior, helping small LLMs learn clinically meaningful decision boundaries from limited data.
+Classical models remain strong on AUROC, but KG-guided LLMs close or surpass this gap while improving macro-AUPR.
+This points toward hybrid designs in which classical or deep risk models provide scores, and KG-guided LLMs act as a clinician-facing layer that produces explanations.
+We also observe encouraging cross-cohort generalization.
+In zero-shot transfer from MIMIC-III to CRADLE, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain stable.
+This pattern indicates that discrimination is preserved, while the induced decision boundaries and reasoning adapt to the target cohort.
 Human evaluation provides complementary evidence that KG guidance improves the clinical usability of explanations.
@@
 Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision.
-ICD–9-to-KG mapping and relevance/path selection rely on GPT-4o and heuristic thresholds, which may introduce subtle errors despite label-consistency filtering.
-Human annotators also highlighted opportunities to improve the input representation itself: our current summaries sometimes blur temporal structure, such as exposure versus outcome windows, and care setting (inpatient versus outpatient), which likely caps attainable performance and can lead even strong models to produce brittle or poorly justified reasoning.
-Finally, the current models require KG access at inference time for KG-conditioned variants.
-Looking ahead, our results point to several concrete directions for future work.
-Applying the framework to richer EHR views, such as notes, medications, labs, and additional KGs could support more nuanced, multimodal reasoning, while replacing simple shortest-path guidance with richer graph reasoning may further improve explanation quality.
-Clinician feedback also points to a complementary direction: improving input structuring by making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and including basic demographics to reduce tendency for the model to hallucinate or over-interpret limited evidence.
-In parallel, future work may introduce temporal or guideline-based constraints to reduce illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
+The mapping of ICD–9 to KG and the selection of the relevance/path rely on GPT-4o and heuristic thresholds, which may introduce subtle errors during data generation.
+Looking ahead, our results suggest several concrete directions.
+Extending the framework to richer EHR views (notes, medications, labs) and additional KGs could support more nuanced multimodal reasoning, while moving beyond simple shortest paths to richer graph reasoning may improve explanation quality.
+In addition, clinician feedback highlights the value of better input structuring—making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and adding basic demographics features (e.g.
+age, sex) —to reduce hallucinations and over-interpretation.
+Future work can introduce temporal or guideline-based constraints to further suppress illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
 Conclusion
@@
 Acknowledgements
-This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University.
-Appendix
-Implementation Details
-All fine-tuning experiments were conducted on a single NVIDIA H200 GPU on our institutional high-performance computing cluster, with 512 GB of host memory allocated via SLURM.
-We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT supervision corpus using full-parameter updates with AdamW (learning rate $1\times 10^{-5}$, cosine learning-rate schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096 tokens, and 10 training epochs in bf16 precision under DeepSpeed ZeRO-3.
+This research was partially supported by internal funds and GPU servers provided by the Computer Science Department of Emory University.
+Implementation Details.
+All fine-tuning experiments were conducted on a single NVIDIA H200 GPU with 512 GB host memory on Emory’s high-performance computing cluster.
+We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT corpus using full-parameter AdamW (learning rate $1\times 10^{-5}$, cosine schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096, and 10 training epochs in bf16 under DeepSpeed ZeRO-3.
 We monitored validation performance and selected the checkpoint with the best macro-AUPR.
 KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100.
-Prompt used:
+The exact prompt templates used in our pipeline are listed below.
 Figure 4: Prompt templates used in our KG-guided CoT pipeline: (top-left) disease-relevant node selection, (top-right) KG path pruning and selection, (bottom) CoT generation conditioned on KG evidence and visit features.

### v3

@@
 Abstract
-Electronic health records (EHRs) enable powerful clinical prediction models but typically offer only coarse, post hoc explanations that are difficult to use in patient-level decision making.
-To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III.
-We map ICD-9 codes to PrimeKG, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes.
-We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models under two resulting training cohorts (400 and 1,000 index visits from MIMIC-III).
-On ten PrimeKG-mapped diseases and small training cohorts, our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47.
-Models transfer zero-shot to the CRADLE cohort without any additional fine-tuning, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
+Electronic health records (EHRs) support strong clinical prediction but often provide coarse, post hoc explanations that are hard to use for patient-level decisions.
+We propose a knowledge-graph (KG)–guided chain-of-thought (CoT) framework for visit-level disease prediction on MIMIC-III.
+We map ICD-9 codes to PrimeKG, mine disease-relevant nodes and paths, and use these paths to scaffold temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes.
+We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models on two small cohorts (400 and 1,000 index visits) across ten PrimeKG-mapped diseases.
+Our models outperform strong classical baselines, reaching AUROC of 0.66–0.70 and macro-AUPR of 0.40–0.47.
+Without additional training, the models transfer zero-shot to the CRADLE cohort, improving accuracy from 0.40–0.51 to 0.72–0.77.
+Blinded clinicians consistently prefer KG-guided CoT for clarity, relevance, and correctness.
+Code is available at https://github.com/JonathanWry/KG-guided-LLM-pipeline.
 Introduction
-Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation; zheng2014automatically As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.choi2016retain; ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
-Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.wei2022chain; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions.
+Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.11; 16; 25 As interest in digital medicine grows,5; 12 these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.10; 6 Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,6 while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.4; 16 Despite these advances in predictive performance, most EHR models provide only coarse, post hoc interpretability (such as global feature importance, attribution scores, or attention weights), offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.6; 4; 1; 17
+Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.20; 19; 9; 21 In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions.
 However, applying CoT in medicine remains challenging.
-High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.xu2024ram; liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability.
-Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
+High-quality, clinically validated CoT datasets are scarce 15; 24 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.23; 13 These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability.
+Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.2; 13
 In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data.
-We propose a KG-guided CoT framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations.
-Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation.
+We propose a KG-guided CoT framework for MIMIC-III10 that produces clinically grounded reasoning rather than unconstrained free-form explanations.
+In our formulation, each visit-level instance uses ICD-9 code features from an index visit $t$ to predict a binary disease label at the immediately subsequent visit t+1t{+}1.
+Concretely, we map ICD-9 codes to PrimeKG2 through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation.
 Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o.
@@
 This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
-We evaluate this framework on ten PrimeKG-mapped diseases in MIMIC-III across small patient cohorts (400 and 1,000 patients), where our KG-anchored CoT supervision enables LLMs to outperform strong classical baselines such as XGBoost and Random Forest, achieving AUROC of approximately 0.67 and 0.70 and macro-AUPR of approximately 0.40 and 0.47, respectively.
+We evaluate this framework on ten PrimeKG-mapped diseases in MIMIC-III across small patient cohorts (400 and 1,000 patient index visits), where our KG-anchored CoT supervision enables LLMs to outperform strong classical baselines such as XGBoost and Random Forest, achieving AUROC of approximately 0.67 and 0.70 and macro-AUPR of approximately 0.40 and 0.47, respectively.
 Clinician evaluations further show that KG-guided CoT yields substantially clearer, more coherent, and more clinically sound reasoning across three dimensions: clarity & coherence, coverage & relevance, and correctness & soundness.
@@
 Method
+Figure 1: General pipeline for KG-guided CoT data generation. Observed clinical features are denoted in Blue; Intermediate KG nodes are denoted in Green, and the Target disease is denoted in Red.
 Preliminaries.
@@
 Since biomedical relations in our external knowledge source are represented as direction- and type-specific (e.g., disease $\rightarrow$ gene and gene $\rightarrow$ disease may be encoded as distinct edges), we treat the underlying biomedical KG as a heterogeneous directed multigraph.
-Specifically, we adopt PrimeKGchandak2023primekg as our external knowledge source for grounding visit-level prognostic reasoning.
+Specifically, we adopt PrimeKG2 as our external knowledge source for grounding visit-level prognostic reasoning.
 PrimeKG is a large-scale biomedical KG that integrates curated relationships among diseases, phenotypes, drugs, genes, and other biomedical entities into a unified graph.
@@
 In our experiments, we instantiate this general setup using ICD–9 diagnosis codes as features, with N=7,423N=7{,}423 and $|\mathcal{D}|=10$ diseases on the MIMIC-III cohort (details in the Results section).
-Figure 1: General pipeline for KG-guided CoT data generation.
 KG Entity Mapping.
@@
 Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{k,d}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$.
-Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
+Following prior work on KG-grounded reasoning 14; 3, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
 $\tilde{\mathcal{P}}_{k,d}=\mathrm{shortest\_path}\!\big(\hat{e}_{k,d},\hat{e}_{d},G,L\big),$
@@
 The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs forms a compact but high-quality KG-grounded supervision corpus.
-We fine-tune lightweight open-weight LLMs (e.g., LLaMA grattafiori2024llama, Gemma team2024gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
+We fine-tune lightweight open-weight LLMs (e.g., LLaMA 7, Gemma 18) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
 Figure 2: Schematic and example of KG-guided CoT generation and filtering.
+Observed clinical features are denoted in Blue; Intermediate KG nodes are denoted in Green, and the Target disease is denoted in Red.
 Results
 Dataset.
-We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits.
+We work with longitudinal visits from the MIMIC-III critical care database.10 Following the setup of Harutyunyan et al.,8 we focus on patients with more than one hospital visit and construct all pairs of adjacent visits.
 For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1.
-This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
+This preprocessing yields 12,353 labeled index visits (i.e., visit pairs).
 Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{N}$ with N=7,423N=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit.
-The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 disease labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1).
+The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).8 In this work, we adopt the same 25 disease labels but evaluate on |D|=10|D|{=}10 diseases that map cleanly to PrimeKG (Table 1).
 For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1.
 Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict.
-We perform visit-level splitting over the 12,353 labeled index visits, randomly reserving 10% of cases as a held-out test set.
-From the remaining pool, we construct training sets of size 400 and 1,000 index visits to study data efficiency, using the rest for model selection.
+e split 12,353 index visits by visit (10% test) and subsample training sets of 400 and 1,000 visits for data-efficiency experiments, using the remainder for validation.
 In addition to MIMIC-III, we evaluate zero-shot transfer on Project CRADLE (Emory Clinical Research Analytics Data Lake Environment), a de-identified EHR repository from Emory Healthcare.
-Following prior work,xu2023hypehr we focus on patients with type 2 diabetes and formulate a binary prediction task: whether a patient will experience a
+Following prior work,22 we focus on patients with type 2 diabetes and formulate a binary prediction task: whether a patient will experience a
 Table 1: The 10 MIMIC-III diseases mapped into PrimeKG and used as prediction targets.
@@
 Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant.
-When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming strong classical baselines such as XGBoost and Random Forest.
+When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming classical baselines such as XGBoost and Random Forest.
 These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
-To evaluate cross-cohort robustness, we apply the MIMIC-trained models directly to the CRADLE dataset (36,611 patients, 12,724 binary features) without any re-training or calibration.
-The prediction task is cardiovascular disease onset within one year of type 2 diabetes diagnosis.
-Table 3: Zero-shot transfer performance on the CRADLE cohort (cardiovascular disease within one year after T2D diagnosis). Best metric in bold, second-best underlined.
+Table 3: Ablation on MIMIC-III visit-level disease prediction with 1000 patient index visits. Best metric in bold, second-best underlined.
+[…]
+LLaMA3-8B + KG (unfiltered) 0.8485 0.6589 0.4473 0.4023
+LLaMA3-8B (no KG, filtered) 0.8613 0.6640 0.4637 0.4357
+[…]
+Table 4: Zero-shot transfer on the CRADLE cohort. Best metric in bold, second-best underlined.
 […]
@@
 […]
-MIMIC-trained models transfer effectively to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes.
+To isolate KG guidance and filtering effects, we compare three variants at 1,000 index visits in Table 3.
+Removing KG guidance or using unfiltered CoT traces reduces AUROC/AUPR, suggesting that ungrounded or noisy rationales degrade ranking performance.
+Overall, KG-grounded and curated CoT supervision is key to strong AUROC/AUPR in this regime.
+To assess cross-cohort robustness, we evaluate the MIMIC-trained models zero-shot on CRADLE (36,611 patients; 12,724 binary features) to predict cardiovascular disease within one year after type 2 diabetes diagnosis, without re-training or calibration.
+MIMIC-trained models transfer to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes.
 For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while modestly raising AUROC (to $\approx 0.52$–0.53) and AUPR (to $\approx 0.25$).
 Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to 0.72–0.77 and macro-F1 from 0.31 to about 0.48, again with modest changes in AUROC and AUPR.
-These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
+These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and spaces, enabling zero-shot transfer of backbones to out-of-distribution EHR data.
 Human Evaluation.
 To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline.
-For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace, one from the KG-guided model and one from the baseline.
-Without knowing which model produced which trace, clinicians made pairwise preferences, selecting the better explanation (A or B) along three dimensions: 1.
-Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions.
-2.
-Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoids unnecessary detail.
-3.
-Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
-For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better.
-The annotation was performed by two human experts from different medical schools: an MD professor and a MD/PhD student in biomedical informatics.
+For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized model outputs (A and B), each containing a prediction and its reasoning trace (one KG-guided, one baseline).
+Without knowing which model produced which trace, clinicians made a pairwise preference judgment (A vs.
+B) along three complementary dimensions:
+  1. 1.
+    Clarity & Coherence (presentation): he trace is easy to follow, logically ordered, and avoids jumps, circular reasoning, or contradictions.
+  2. 2.
+    Coverage & Relevance (evidence selection): the trace surfaces key patient-specific factors for the target disease, stays clinically focused, and avoids generic or distracting content.
+  3. 3.
+    Correctness & Soundness (clinical validity): clinical claims and timelines are medically plausible and consistent with the provided features/outcome (no fabricated details or inappropriate causal leaps), and the conclusion follows from the stated evidence.
+These dimensions are intentionally distinct: clarity evaluates how the reasoning is communicated, coverage evaluates what evidence is selected and emphasized, and correctness evaluates whether the medical content is accurate and logically warranted.
+The annotation was performed by two human experts from different medical schools: an MD professor and an MD/PhD student in biomedical informatics.
 In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the MD/PhD student reviewing 50 cases.
-Figure 3: Clinician preferences for reasoning quality.
-On the 115 annotated cases, both clinicians preferred the KG-guided model (Model A) over the untuned baseline (Model B), with overall 96.5% of cases for Clarity & Coherence, 94.8% of cases for Coverage & Relevance, and 98.3% of cases for Correctness & Soundness.
-Qualitatively, clinicians reported that KG-guided CoT traces more often (i) organized evidence in a temporally consistent manner, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details.
-By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES even when only weak signals were present).
-While both models are sensitive to poor or incomplete input structure, the untuned baseline was more prone to repetitive differentials and illogical reasoning, especially on shorter or noisier inputs.
+Inter-rater consistency.
+To quantify consistency across adjudicators, we additionally double-annotated a random subset of $n_{\text{overlap}}$ of 55 cases (both clinicians independently rated the same cases across all three dimensions) and computed Cohen’s $\kappa$ on the binary preference decision (KG-guided vs.
+baseline).
+Agreement was $\kappa_{\text{clar}}$ for Clarity & Coherence, $\kappa_{\text{cov}}$ for Coverage & Relevance, and $\kappa_{\text{corr}}$ for Correctness & Soundness.
+Figure 3: Clinician preference for KG-guided explanations vs. the untuned baseline across three dimensions (donuts show % preferred; remainder is baseline).
+On the 115 annotated cases, both clinicians preferred the KG-guided model (Model A) over the untuned baseline (Model B): 96.5% for Clarity & Coherence, 94.8% for Coverage & Relevance, and 98.3% for Correctness & Soundness.
+Qualitatively, clinicians noted that KG-guided CoT traces more often (i) organized evidence temporally, (ii) highlighted clinically meaningful risk factors over superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details.
+By contrast, the untuned baseline was described as more “trigger-happy” (more likely to conclude YES from weak signals).
+While both models are sensitive to poor or incomplete inputs, the baseline was more prone to repetitive differentials and illogical reasoning, especially on short or noisy inputs.
 Discussion
-In this study, we proposed a KG-guided chain-of-thought generation framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data.
-By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we construct a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs.
-On MIMIC-III visit-level disease prediction, KG-guided CoT improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and makes them competitive with strong classical baselines.
-With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$; with 1,000 visits, performance increases to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$.
-These results suggest that KG-anchored CoT supervision acts as an effective reasoning prior, helping small LLMs learn clinically meaningful decision boundaries from limited data.
-Classical models remain strong on AUROC, but KG-guided LLMs close or surpass this gap while improving macro-AUPR.
-This points toward hybrid designs in which classical or deep risk models provide scores, and KG-guided LLMs act as a clinician-facing layer that produces explanations.
-We also observe encouraging cross-cohort generalization.
-In zero-shot transfer from MIMIC-III to CRADLE, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain stable.
-This pattern indicates that discrimination is preserved, while the induced decision boundaries and reasoning adapt to the target cohort.
-Human evaluation provides complementary evidence that KG guidance improves the clinical usability of explanations.
-In 115 blinded case comparisons between our KG-guided LLaMA model and an untuned LLM baseline, the KG-guided traces were preferred on clarity and coherence, coverage and relevance, and correctness and soundness in the large majority of cases.
-Qualitative feedback from the MD professor and PhD student indicated that KG-guided traces more often (i) organized evidence in a temporally consistent way, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details.
-By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES) and sometimes reached correct labels via illogical or poorly justified reasoning, especially on shorter or noisier inputs.
-These observations suggest that anchoring CoT to KG paths, combined with label-consistency filtering, shapes both the style and the content of reasoning toward disease-specific, clinically grounded narratives.
-Several limitations remain.
-Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision.
-The mapping of ICD–9 to KG and the selection of the relevance/path rely on GPT-4o and heuristic thresholds, which may introduce subtle errors during data generation.
-Looking ahead, our results suggest several concrete directions.
-Extending the framework to richer EHR views (notes, medications, labs) and additional KGs could support more nuanced multimodal reasoning, while moving beyond simple shortest paths to richer graph reasoning may improve explanation quality.
-In addition, clinician feedback highlights the value of better input structuring—making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and adding basic demographics features (e.g.
-age, sex) —to reduce hallucinations and over-interpretation.
-Future work can introduce temporal or guideline-based constraints to further suppress illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
+We introduce a KG-guided chain-of-thought (CoT) framework that uses PrimeKG structure to scaffold LLM reasoning over structured EHRs.
+By mapping ICD–9 codes to PrimeKG,2 mining disease-specific relevance sets and short paths, and applying label-consistency filtering, we build a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs.
+Finding 1: KG-anchored CoT improves low-data learning, especially under imbalance.
+Across both 400 and 1,000 labeled index visits (Table Experiment:), KG-guided CoT substantially improves LLaMA-3.1-8B and Gemma-7B and is competitive with strong classical baselines.
+Gains are largest on macro-AUPR (e.g., $\approx 0.25\rightarrow 0.40$ for Gemma at 400), suggesting that KG-anchored supervision encourages disease-relevant evidence over superficial code co-occurrence.
+Finding 2: Filtering and KG evidence play complementary roles.
+In the ablation (Table 3), filtering alone yields the best accuracy/F1, while adding KG evidence yields the best AUROC (up to $\approx 0.70$).
+Practically, filtering reduces noisy supervision and spurious positives, whereas KG structure strengthens ranking; threshold calibration can further improve accuracy/F1.
+Future work can further improve threshold-dependent metrics (accuracy/F1) by calibrating decision thresholds per disease or cohort.
+Finding 3: Zero-shot transfer mainly shifts the operating point, motivating calibration.
+On CRADLE (Table 4), fine-tuned models boost accuracy (0.72–0.77) with only modest AUROC/AUPR changes, indicating a changed decision rule under a fixed threshold rather than large ranking gains.
+Deployment evaluations should therefore include calibration and threshold selection, especially across cohorts with different prevalences and coding practices.
+Finding 4: Clinicians preferred KG-guided explanations.
+Clinicians preferred KG-guided traces on clarity/coherence, coverage/relevance, and correctness/soundness for most cases.
+They cited better temporal organization, more clinically meaningful risk factors, and fewer unsupported leaps, suggesting KG evidence as an effective planning scaffold for clinician-facing rationales.
+Limitations and future work.
+Our approach depends on the coverage and quality of PrimeKG, so gaps or biases in the KG can propagate into both supervision and explanations.
+The ICD to KG mapping and the relevance/path selection steps rely on GPT-4o and heuristic constraints.
+While we constrain candidate paths using shortest-path enumeration and use the LLM only as a selector, the pruned set is not guaranteed to be exhaustive: alternative clinically valid paths may be removed in favor of a compact scaffold.
+Methodologically, we also adopt a single-visit formulation ($x_{t}$) for next-visit prediction, which simplifies evaluation but does not fully exploit longitudinal history.
+Future work can extend the framework to richer EHR views (notes, medications, labs), incorporate multi-visit context, and integrate clinician-in-the-loop validation for KG evidence selection and pruning.
 Conclusion
@@
 Acknowledgements
-This research was partially supported by internal funds and GPU servers provided by the Computer Science Department of Emory University.
-Implementation Details.
-All fine-tuning experiments were conducted on a single NVIDIA H200 GPU with 512 GB host memory on Emory’s high-performance computing cluster.
-We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT corpus using full-parameter AdamW (learning rate $1\times 10^{-5}$, cosine schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096, and 10 training epochs in bf16 under DeepSpeed ZeRO-3.
-We monitored validation performance and selected the checkpoint with the best macro-AUPR.
-KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100.
-The exact prompt templates used in our pipeline are listed below.
-Figure 4: Prompt templates used in our KG-guided CoT pipeline: (top-left) disease-relevant node selection, (top-right) KG path pruning and selection, (bottom) CoT generation conditioned on KG evidence and visit features.
+This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University, the US National Science Foundation under Award Numbers 2442172, 2312502, 2319449, and the US National Institutes of Health under Award Numbers K25DK135913, RF1NS139325, R01DK143456 and U18DP006922.